# Holomorphic Deformations of Hyperbolicity Notions on Compact Complex Manifolds

### v1

@@ A Hermitian metric $\omega$ on $X$ is said to be :
    A $(p,p)$-form $u\in\Lambda^{p,p}(E^{*})$ is tranverse if for any linearly independent $(1,0)$-forms $\alpha_{1},\dots,\alpha_{q}\in E^{*}$ we have :
@@
    Let $X$ be a compact complex manifold with $\dim_{\mathbb{C}}X=n$. A $(p,p)$-form $\alpha\in\mathscr{C}^{\infty}_{p,p}(X)$ is transverse if $\forall x\in X$, $\alpha(x)\in\Lambda^{p,p}T_{x}^{*}X$ is transverse in the sense of $(1)$.
@@
All transverse $(p,p)$-forms $u$ are real.
Here, we recall the generalizations of the special metrics in degree $1\leq p\leq n-1$.
@@
    p-Kähler if it admits a $d$-closed transverse $(p,p)$-form $\Omega$. Here, $\Omega$ is called a p-Kähler form ((Definition 1.11, [3])).
@@
    p-Hermitian symplectic ((or p-HS for short)) if it admits a real $d$-closed $2p$-form $\Omega$ such that $\Omega^{p,p}$ is transverse. Here, $\Omega^{p,p}$ is called a p-HS form ((Definition 1.11, [3])).
@@
    p-SKT ((or p-pluriclosed)) if it admits a $\partial\bar{\partial}$-closed transverse $(p,p)$-form $\Omega$. Here, $\Omega$ is called a p-SKT form ((Theorem 2.4, [1])).
@@
    p-HS hyperbolic if it admits a d~($\widetilde${d}(bounded)) $p$-HS form.
@@
    One can define astheno-Kähler hyperbolicity in the same vein : We say that a compact astheno-Kähler manifold $X$ is astheno-Kähler hyperbolic if it admits an astheno-Kähler metric $\omega$ such that $\omega^{n-2}$ is ($\widetilde{\partial+\bar{\partial}}$)(bounded). Then :
@@
    By arguments similar to those used in Propositions 3.4.(3) and 3.6, one can show that a Kähler hyperbolic manifold is $p$-Kähler hyperbolic, and every HS hyperbolic manifold is $p$-HS hyperbolic, respectively, for any $1\leq p\leq n-1$.
@@
Let $M$ be a complete simply connected Riemannian manifold with sectional curvature $K\leq-c^{2}<0$. Then every closed bounded $k$-form on $M$ is d((bounded)), for any $k\geq 2$.
@@
Let $X$ be a compact $p$-Kähler ((resp. $p$-HS)) manifold with $\dim_{\mathbb{C}}X=n$, homotopic to a compact negatively curved Riemannian manifold $(Y,g_{Y})$, with $1\leq p\leq n-1$. Then $X$ is $p$-Kähler hyperbolic ((resp. $p$-HS hyperbolic)).
@@
We will carry out the proof in the $p$-HS case, the $p$-Kähler case can be handled in a similar manner. Let $\Omega$ be a real $d$-closed $2p$-form on $X$ such that $\Omega^{p,p}$ is transverse, i.e. $\Omega^{p,p}$ is a $p$-HS form. The goal is to show that $\Omega$ is $\widetilde{d}$(bounded). Let $\pi_{X}:\widetilde{X}\longrightarrow X$ and $\pi_{Y}:\widetilde{Y}\longrightarrow Y$ be the universal coverings of $X$ and $Y$ respectively. Since $X$ and $Y$ are homotopic, there exist two smooth maps $f_{1}:X\longrightarrow Y$ and $f_{2}:Y\longrightarrow X$ such that the induced map $(f_{2}\circ f_{1})^{*}$ on $H^{*}_{DR}(X,\mathbb{R})$ is the identity. Thus for any $\Omega^{\prime}\in[\Omega]\in H^{2p}_{DR}(X,\mathbb{R})$, there exists a (2p12p-1)-form $u$ in $X$ such that :
@@
The case $p=1$ in Theorem 3.14 is already of particular interest. Indeed, if we consider a compact Hermitian symplectic manifold $X$ with negative sectional curvature, then $X$ is HS hyperbolic. However, if we take a Hermitian metric $\omega$ on $X$ (not necessarily making $X$ Hermitian symplectic) with negative sectional curvature (for the Chern connection), then the $(1,1)$-form :
$\eta:=-\mathrm{Ric}^{(1)}(\omega)\underset{\mathrm{loc}}{=}i\partial\bar{\partial}\log(\omega^{n})$ (3.6)
defines a Kähler form on $X$. Hence, $X$ is even Kähler hyperbolic. Therefore, finding an example of a compact non-Kähler, Hermitian symplectic manifold that is homotopic to a compact Riemannian manifold with negative sectional curvature would amount to providing an example of a non-Kähler Kobayashi hyperbolic manifold.
@@
An interesting question is whether an analogue of Theorem 3.14 holds in the $p$-SKT hyperbolic setting (maybe with the sectional curvature replaced by the holomorphic sectional (or bisectional) curvature ?).
@@
    If $Y$ is $p$-HS ((resp. $p$-HS hyperbolic)) for all $q\leq p<m$, $X\times Y$ is $(p+n)$-HS ((resp. $(p+n)$-HS hyperbolic)) for all $q\leq p<m$.
@@
    If $Y$ is $p$-SKT ((resp. $p$-SKT hyperbolic)) for all $q\leq p<m$, $X\times Y$ is $(p+n)$-SKT ((resp. $(p+n)$-SKT hyperbolic)) for all $q\leq p<m$.
@@
    For $q\leq p<n$, we let $\Omega_{p}$ be a real $d$-closed $2p$-form on $Y$ such that $\Omega^{p,p}$ is transverse, i.e. $\Omega^{p,p}$ is a $p$-HS form. Fix $q+n\leq j<m+n$ and let $h=\max\{0,j-m\}$. We set :
@@
    is transverse. Hence $\Theta_{j}^{j,j}$ defines a $j$-HS form on $X\times Y$. If moreover, we assume the $p$-forms $\Omega_{p}$ are $\widetilde{d}$(bounded), i.e. $\forall q\leq p<m$, there exists a bounded (2p12p-1)-form $\eta_{p}$ in $\widetilde{Y}$ such that $\pi_{Y}^{*}\Omega_{p}=d\eta_{p}$. By pulling-back the expression (3.7) to $\widetilde{X}\times\widetilde{Y}$, we get :
@@
We now aim to highlight the relationship between the hyperbolicity notions defined in degree $p$ and their corresponding geometric counterparts. Recall that in the case $p=1$, these notions are those of Kähler hyperbolicity, HS hyperbolicity, and SKT hyperbolicity (Remark 3.12.1), which have been shown to imply Kobayashi hyperbolicity (see [11], Theorem 4.1, Corollary 3.5 and [36], Theorem 3.5, respectively).
For p=n1p=n-1, the corresponding notions are Balanced hyperbolicity, sG hyperbolicity and Gauduchon hyperbolicity (Remark 3.12.2), which have been proven to be divisorially hyperbolic (see [38], Theorem 2.8; [35], Theorem 2.6; and [36], Theorem 3.10, respectively).
@@
Let $(X,\omega)$ be a compact Hermitian manifold with $\dim_{\mathbb{C}}X=n\geq 2$, and let $f:\mathbb{C}^{p}\longrightarrow X$ be a holomorphic map with $1\leq p\leq n-1$, which is non-degenerate at some point $x_{0}\in\mathbb{C}^{p}$, i.e. $d_{x_{0}}f:\mathbb{C}^{p}\longrightarrow T^{1,0}_{x_{0}}X$ is of maximal rank. Set : $\Sigma_{f}:=\{x\in\mathbb{C}^{p}\hskip 1.0pt:\hskip 1.0ptf\text{ is degenerate at }x\}$; this is a proper analytic subset of $\mathbb{C}^{p}$. Then $f^{*}\omega$ defines a degenerate metric on $\mathbb{C}^{p}$ with degeneration set $\Sigma_{f}$ ; that is $f^{*}\omega\geq 0$ on $\mathbb{C}^{p}$ and defines a genuine Hermitian metric on $\mathbb{C}^{p}\setminus\Sigma_{f}$. For any $r>0$, we define the ($\omega,f$)-volume of the ball $\mathbb{B}_{r}\subset\mathbb{C}^{p}$ of radius $r$ centred at 0 to be :
@@
This implies that the (2p12p-1)-form $d\sigma_{\omega,f}:=*_{f^{*}\omega}\left(\dfrac{d\beta(z)}{\lvert d\beta(z)\rvert_{f^{*}\omega}}\right)$ defines an area measure on spheres $\mathbb{S}_{r}:\partial\mathbb{B}_{r}\subset\mathbb{C}^{p}$. In particular, we define the ($\omega,f$)-area of the sphere $\mathbb{S}_{r}$ of radius $r>0$ by :
@@
Within the same setting, we say that $f$ has subexponential growth if :
@@
    When $p=1$, it is clear that Kobayashi hyperbolicity implies $1$-cyclic hyperbolicity. Conversly, assume $X$ is a compact complex manifold which admits a non-constant holomorphic map $f:\mathbb{C}\longrightarrow X$. Thanks to Brody’s Renormalisation Lemma ([9], Lemma 2.1), we can assume $f$ of order $2$ (see [33], Theorem 2.6-p.72), that is :
@@
    which is clearly dominated by $r\text{Vol}_{\omega,f}(\mathbb{B}_{r})$ for $r>r_{0}$ large enough. So $f$ satisfies also condition (i) in the Definition of subexponential growth. Thus, $1$-cyclic hyperbolicity coincides with Kobayashi hyperbolicity.
@@
S. Marouani and F. Haggui established that every $p$-Kähler hyperbolic manifold is $p$-cyclically hyperbolic (Theorem 3.15 in [25]). Regarding $p$-HS hyperbolicity, M. Yi showed that this condition also implies $p$-cyclic hyperbolicity (Theorem 3.2 in [35], where $p$-HS is referred to as weakly $p$-Kähler). We now prove that the same holds for $p$-SKT hyperbolic manifolds. Before doing so, we first record the following observation :
Observation 3.22.
Let ($X,\omega$) be a compact Hermitian manifold with $\dim_{\mathbb{C}}X=n$ and $\pi:\widetilde{X}\longrightarrow X$ be the universal cover of $X$, and let $f:\mathbb{C}^{p}\longrightarrow X$ a holomorphic map that is non-degenerate at some point $x_{0}\in\mathbb{C}^{p}$, with $1\leq p\leq n-1$. Then $f^{*}\omega$ defines a degenerate Hermitian metric on $X$ (with degeneracy locus $\Sigma_{f}$). Since $\mathbb{C}^{p}$ is simply connected, $f$ lifts to a holomorphic map $\widetilde{f}:\mathbb{C}^{p}\longrightarrow\widetilde{X}$, that is : $\pi\circ\widetilde{f}=f$. In particular, the degeneracy locus of $\widetilde{f}$ is also $\Sigma_{f}$. Let $\eta$ be a $\widetilde{\omega}-$bounded $k$-form on $\widetilde{X}$, with $k\leq 2p$. For any $x\in\mathbb{C}^{p}$, and any tangent vectors $\xi_{1},\cdots,\xi_{k}\in\mathbb{C}^{p}$ at $x$, we have :
@@
Hence $\widetilde{f}^{*}\eta$ is $f^{*}\omega$-bounded on $\mathbb{C}^{p}$.
Theorem 3.23.
Proof.
Let ($X,\omega$) be a compact Hermitian manifold with $\dim_{\mathbb{C}}X=n$, equipped with a $p$-SKT hyperbolic form $\Omega$, i.e. $\Omega$ is a smooth transverse ($p,p$)-form such that $\partial\bar{\partial}\Omega=0$ and $\widetilde{\Omega}=\partial\alpha+\bar{\partial}\gamma$, where $\alpha$ (resp. $\gamma$) is a smooth $\widetilde{\omega}-$bounded (p1,pp-1,p)-form (resp. (p,p1p,p-1)-form) on $\widetilde{X}$. Assume that a holomorphic map $f:\mathbb{C}^{p}\longrightarrow X$ such as in Definition 3.18. Then $f^{*}\Omega$ defines a positive ($p,p$)-form (hence strongly positive by maximality of the dimension, see Corollary 1.9 in [13], §.1, Chapter III) on $\mathbb{C}^{p}\setminus\Sigma_{f}$. Hence, there exists a smooth positive function $h:\mathbb{C}^{p}\setminus\Sigma_{f}\longrightarrow\mathbb{R}$ such that : $f^{*}\omega_{p}=hf^{*}\Omega$, where $\omega_{p}=\dfrac{\omega^{p}}{p!}$. In particular, we have for any $z\in\mathbb{C}^{p}\setminus\Sigma_{f}$ :
$\sqrt{\binom{n}{p}}=\lvert\omega_{p}(f(z))\rvert_{\omega}=\lvert f^{*}\omega_{p}(z)\rvert_{f^{*}\omega}=h(z)\lvert f^{*}\Omega(z)\rvert_{f^{*}\omega}=h\lvert\Omega(f(z))\rvert_{\omega}$
Thus : $h(z)=\displaystyle\sqrt{\binom{n}{p}}\lvert\Omega(f(z))\rvert^{-1}_{\omega}$, $\forall z\in\mathbb{C}^{p}\setminus\Sigma_{f}$. On the other hand, since $\Omega(x)\neq 0$ for any $x\in X$ by transversailty, the compactness of $X$ implies that there exists a constant $A>0$ such that : $\dfrac{1}{A}\leq\lvert\Omega(x)\rvert_{\omega}\leq A$. Hence, we get for any $r>0$ :
$\text{Vol}_{\omega,f}(\mathbb{B}_{r})=\int_{\mathbb{B}_{r}}f^{*}\omega_{p}\overset{(a)}{=}\int_{\mathbb{B}_{r}}hf^{*}\Omega\leq A\sqrt{\binom{n}{p}}\int_{\mathbb{B}_{r}}f^{*}\Omega$ (3.15)
where ($a$) is due to the fact that $\Sigma_{f}$ has zero measure. Moreover, we have :
$f^{*}\Omega=\widetilde{f}^{*}\widetilde{\Omega}=\widetilde{f}(\partial\alpha+\bar{\partial}\gamma)\overset{(b)}{=}\widetilde{f}(d(\alpha+\gamma))=d\widetilde{f}(\alpha+\gamma)$ (3.16)
where ($b$) is a consequence of the bidegree constraints. Substituting (3.16) into (3.15) gives :
$\text{Vol}_{\omega,f}(\mathbb{B}_{r})\leq A\sqrt{\binom{n}{p}}\int_{\mathbb{B}_{r}}d\widetilde{f}(\alpha+\gamma)=A\sqrt{\binom{n}{p}}\int_{\mathbb{S}_{r}}\widetilde{f}(\alpha+\gamma)\overset{(c)}{\leq}BA_{\omega,f}(\mathbb{S}_{r})$ (3.17)
Where ($c$) follows from Observation 3.22 (here $\eta=\alpha+\gamma$). On the other hand, the Cauchy-Schwarz inequality yields :

### v4

@@ 3.2  $p$-SKT and $p$-HS Hyperbolicity
    A $(p,p)$-form $u\in\Lambda^{p,p}(E^{*})$ is weakly strictly positive if for any linearly independent $(1,0)$-forms $\alpha_{1},\dots,\alpha_{q}\in E^{*}$ we have:
@@
    Let $X$ be a compact complex manifold with $\dim_{\mathbb{C}}X=n$. A $(p,p)$-form $\alpha\in\mathscr{C}^{\infty}_{p,p}(X)$ is weakly strictly positive if $\forall x\in X$, $\alpha(x)\in\Lambda^{p,p}T_{x}^{*}X$ is weakly strictly positive in the sense of $(1)$.
@@
All weakly strictly positive $(p,p)$-forms $u$ are real, i.e. $\bar{u}=u$.
Here, we recall the generalizations of special metrics in degree $2p$, with $1\leq p\leq n-1$.
@@
    p-Kähler if it admits a $d$-closed weakly strictly positive $(p,p)$-form $\Omega$. Here, $\Omega$ is called a p-Kähler form ((Definition 1.11, [3])).
@@
    p-Hermitian-symplectic ((or p-HS for short)) if it admits a real $d$-closed $2p$-form $\Omega$ such that $\Omega^{p,p}$ is weakly strictly positive. Here, $\Omega^{p,p}$ is called a p-HS form ((Definition 1.11, [3])).
@@
    p-SKT ((or p-pluriclosed)) if it admits a $\partial\bar{\partial}$-closed weakly strictly positive $(p,p)$-form $\Omega$. Here, $\Omega$ is called a p-SKT form ((Theorem 2.4, [1])).
@@
    p-HS hyperbolic if it admits a d~($\widetilde${d}(bounded)) $p$-HS form ((this was referred to as weakly $p$-Kähler hyperbolic in Definition 3.1, [40])).
@@
    One can define astheno-Kähler hyperbolicity in a similar way: We say that a compact astheno-Kähler manifold $X$ is astheno-Kähler hyperbolic if it admits an astheno-Kähler metric $\omega$ such that $\omega^{n-2}$ is ($\widetilde{\partial+\bar{\partial}}$)(bounded). Then:
@@
    By arguments similar to those used in Propositions 3.6.(3) and 3.8, one can show that a Kähler hyperbolic manifold is $p$-Kähler hyperbolic, and every HS hyperbolic manifold is $p$-HS hyperbolic, respectively, for any $1\leq p\leq n-1$.
@@
Let $M$ be a complete simply connected Riemannian manifold with Riemannian sectional curvature $K\leq-c^{2}<0$. Then every closed bounded $k$-form on $M$ is d((bounded)), for any $k\geq 2$.
Note that such a manifold $M$ cannot be compact since the Cartan–Hadamard theorem states that a complete simply connected Riemannian manifold with non-positive sectional curvature is diffeomorphic to $\mathbb{R}^{\dim M}$ via the exponential map at any point.
@@
Let $X$ be a compact $p$-Kähler ((resp. $p$-HS)) manifold with $\dim_{\mathbb{C}}X=n$, homotopic to a compact Riemannian manifold $(Y,g_{Y})$ with negative Riemannian sectional curvature, with $1\leq p\leq n-1$. Then $X$ is $p$-Kähler hyperbolic ((resp. $p$-HS hyperbolic)).
@@
We will spell out the proof in the $p$-HS case, the $p$-Kähler case can be handled in a similar manner. Let $\Omega$ be a real $d$-closed $2p$-form on $X$ such that $\Omega^{p,p}$ is weakly strictly positive, i.e. $\Omega^{p,p}$ is a $p$-HS form. The goal is to show that $\Omega$ is $\widetilde{d}$(bounded). Let $\pi_{X}:\widetilde{X}\longrightarrow X$ and $\pi_{Y}:\widetilde{Y}\longrightarrow Y$ be the universal coverings of $X$ and $Y$ respectively. Since $X$ and $Y$ are homotopic, there exist two smooth maps $f_{1}:X\longrightarrow Y$ and $f_{2}:Y\longrightarrow X$ such that the induced map $(f_{2}\circ f_{1})^{*}$ on $H^{*}_{DR}(X,\mathbb{R})$ is the identity. Thus for any $\Omega^{\prime}\in[\Omega]\in H^{2p}_{DR}(X,\mathbb{R})$, there exists a (2p12p-1)-form $u$ in $X$ such that:
@@
  1. 1.
    Theorem 3.16 is not relevant when the manifold $X$ itself admits a Hermitian metric $\omega$ with negative sectional curvature (for the Chern connection), since this implies the negativity of the first Chern-Ricci curvature form of $\omega$. Then, the ($1,1$)-form:
    $\eta:=-\mathrm{Ric}^{(1)}(\omega)=i\Theta_{\omega}(K_{X})\underset{\mathrm{loc}}{=}i\partial\bar{\partial}\log(\omega^{n}).$
    defines a Kähler form on $X$. In particular, $X$ is Kähler hyperbolic since ($X,\omega$) is diffeomorphic to ($X,\eta$) and are therefore homotopic. Hence Theorem 3.16 applies.
  2. 2.
    The case $p=1$ in Theorem 3.16 is already of particular interest. Indeed, constructing a compact non-Kähler Hermitian–symplectic manifold that is homotopic to a compact Riemannian manifold with $K<0$ would provide an example of a non-Kähler Kobayashi hyperbolic manifold, assuming an affirmative answer to the question raised by J. Streets and G. Tian in Question 1.7 of [60] concerning the existence of compact non-Kähler Hermitian–symplectic manifolds. Such a scenario is, of course, widely regarded as unlikely in view of the Kobayashi conjecture, which predicts the ampleness of the canonical bundle $K_{X}$ for any compact Kobayashi hyperbolic manifold $X$. Nevertheless, the existence of a compact non-Kähler Kobayashi hyperbolic manifold would simply indicate that the Kähler assumption in the Kobayashi conjecture must be made explicit.
@@
Motivated by the search for non-Kähler Kobayashi hyperbolic manifolds, we ask whether Theorem 1.1 in [9] can be strengthened in the non-Kähler setting: that is, whether a compact complex manifold carrying an SKT metric with negative holomorphic sectional curvature ($\text{HSC}<0$) must be SKT hyperbolic. Recall that any compact Hermitian manifold with $\text{HSC}<0$ is Kobayashi hyperbolic by [34], Theorem 4.1. As recently pointed out to the author by J. Streets, the negativity of the real bisectional curvature already forces a compact Hermitian manifold to be Kähler with ample canonical bundle (see [37], Theorem 1.1), indicating that even slight strengthening of the $\text{HSC}<0$" assumption may preclude genuinely non-Kähler examples. Note, for example, that a compact complex manifold with negative holomorphic bisectional curvature ($\mathrm{HBC}<0$) need not be Kähler SKT hyperbolic. Indeed, manifolds of the latter type have infinite fundamental group, since they are $L^{2}$-Kähler hyperbolic and therefore have generically large fundamental group by Corollary 2.12 in [33]. In contrast, there exist simply connected Kähler manifolds with $\mathrm{HBC}<0$, arising for instance as complete intersections in complex projective manifolds [46].
In [32], Section §.5, it was observed that one does not need to assume negativity of the sectional curvature; instead, a weaker curvature-type condition introduced in the context of pluriclosed star split manifolds [57] suffices to obtain hyperbolicity results in several settings. This suggests the possibility of exploring whether analogous conditions could imply other forms of hyperbolicity for compact complex manifolds, possibly in a partial sense. For instance, one may ask whether such curvature-type hypotheses could lead to a partial hyperbolicity of Hermitian–Symplectic, SKT, strongly Gauduchon, or Gauduchon type. Since partial hyperbolicity is weaker and more flexible than full Kobayashi hyperbolicity, it could potentially allow for a wider range of non-Kähler examples, opening the door to new constructions and examples in this direction.
@@
    If $Y$ is $p$-HS ((resp. $p$-HS hyperbolic)) for all $q\leq p<m$, then $X\times Y$ is $(p+n)$-HS ((resp. $(p+n)$-HS hyperbolic)) for all $q\leq p<m$.
@@
    If $Y$ is $p$-SKT ((resp. $p$-SKT hyperbolic)) for all $q\leq p<m$, then $X\times Y$ is $(p+n)$-SKT ((resp. $(p+n)$-SKT hyperbolic)) for all $q\leq p<m$.
@@
    For $q\leq p<n$, we let $\Omega_{p}$ be a real $d$-closed $2p$-form on $Y$ such that $\Omega^{p,p}$ is weakly strictly positive, i.e. $\Omega^{p,p}$ is a $p$-HS form. Fix $q+n\leq j<m+n$ and let $h=\max\{0,j-m\}$. We set:
@@
    is weakly strictly positive. Hence $\Theta_{j}^{j,j}$ defines a $j$-HS form on $X\times Y$. If moreover, we assume the $p$-forms $\Omega_{p}$ are $\widetilde{d}$(bounded), i.e. $\forall q\leq p<m$, there exists a bounded (2p12p-1)-form $\eta_{p}$ in $\widetilde{Y}$ such that $\pi_{Y}^{*}\Omega_{p}=d\eta_{p}$. By pulling-back the expression (3.6) to $\widetilde{X}\times\widetilde{Y}$, we get:
@@
We now aim to highlight the relationship between the hyperbolicity notions defined in degree $2p$ and their corresponding geometric counterparts. Recall that in the case $p=1$, these are the notions of Kähler hyperbolicity, HS hyperbolicity, and SKT hyperbolicity (Remark 3.14.1). They have been shown to imply Kobayashi hyperbolicity (see [12], Theorem 4.1, Corollary 3.7 and [41], Theorem 3.5, respectively).
For p=n1p=n-1, the corresponding notions are balanced hyperbolicity, sG hyperbolicity and Gauduchon hyperbolicity (Remark 3.14.2). These manifolds have been proven to be divisorially hyperbolic (see [43], Theorem 2.8; [40], Theorem 2.6; and [41], Theorem 3.10, respectively).
@@
Let $(X,\omega)$ be a compact Hermitian manifold with $\dim_{\mathbb{C}}X=n\geq 2$, and let $f:\mathbb{C}^{p}\longrightarrow X$ be a holomorphic map with $1\leq p\leq n-1$, which is non-degenerate at some point $x_{0}\in\mathbb{C}^{p}$, i.e. $d_{x_{0}}f:\mathbb{C}^{p}\longrightarrow T^{1,0}_{x_{0}}X$ is of maximal rank. Set: $\Sigma_{f}:=\{x\in\mathbb{C}^{p}\hskip 1.0pt:\hskip 1.0ptf\text{ is degenerate at }x\}$; this is a proper analytic subset of $\mathbb{C}^{p}$. Then $f^{*}\omega$ defines a degenerate metric on $\mathbb{C}^{p}$ with degeneration set $\Sigma_{f}$ ; that is $f^{*}\omega\geq 0$ on $\mathbb{C}^{p}$ and defines a genuine Hermitian metric on $\mathbb{C}^{p}\setminus\Sigma_{f}$. For any $r>0$, we define the ($\omega,f$)-volume of the ball $\mathbb{B}_{r}\subset\mathbb{C}^{p}$ of radius $r$ centered at 0 to be:
@@
This implies that the (2p12p-1)-form $d\sigma_{\omega,f}:=*_{f^{*}\omega}\left(\dfrac{d\beta(z)}{\lvert d\beta(z)\rvert_{f^{*}\omega}}\right)$ defines an area measure on spheres $\mathbb{S}_{r}:\partial\mathbb{B}_{r}\subset\mathbb{C}^{p}$.
In particular, we define the ($\omega,f$)-area of the sphere $\mathbb{S}_{r}$ of radius $r>0$ by:
@@
In the same setting, we say that $f$ has subexponential growth if:
@@
    When $p=1$, it is clear that Kobayashi hyperbolicity implies $1$-cyclic hyperbolicity. Conversly, assume $X$ is a compact complex manifold which admits a non-constant holomorphic map $f:\mathbb{C}\longrightarrow X$. Thanks to Brody’s Renormalisation Lemma ([10], Lemma 2.1), we can assume $f$ to be of order $2$ (see [36], Theorem 2.6-p.72), that is:
@@
    which is clearly dominated by $r\text{Vol}_{\omega,f}(\mathbb{B}_{r})$ for $r>r_{0}$ large enough. So $f$ satisfies also condition (i) in Definition 3.20. Thus, $1$-cyclic hyperbolicity coincides with Brody hyperbolicity (equivalently, with Kobayashi hyperbolicity).
@@
S. Marouani and F. Haggui established that every $p$-Kähler hyperbolic manifold is $p$-cyclically hyperbolic (Theorem 3.15 in [27]). Regarding $p$-HS hyperbolicity, Y. Ma showed that this condition also implies $p$-cyclic hyperbolicity in [40], Theorem 3.2. We now prove that the same holds for $p$-SKT hyperbolic manifolds:
Theorem 3.24.
Proof.
Let ($X,\omega$) be a compact Hermitian manifold with $\dim_{\mathbb{C}}X=n$ and $\pi:\widetilde{X}\longrightarrow X$ be the universal cover of $X$. Assume $X$ equipped with a $p$-SKT hyperbolic form $\Omega$, i.e. $\Omega$ is a smooth weakly strictly positive ($p,p$)-form such that $\partial\bar{\partial}\Omega=0$ and $\widetilde{\Omega}=\partial\alpha+\bar{\partial}\gamma$, where $\alpha$ (resp. $\gamma$) is a smooth $\widetilde{\omega}-$bounded (p1,pp-1,p)-form (resp. (p,p1p,p-1)-form) on $\widetilde{X}$. Assume that a holomorphic map $f:\mathbb{C}^{p}\longrightarrow X$ such as in Definition 3.20 exists. Then $f^{*}\omega$ defines a degenerate Hermitian metric on $X$ (with degeneracy locus $\Sigma_{f}$). Let us first make the following observation:
Observation 3.25.
Let $\eta$ be an $\widetilde{\omega}-$bounded $k$-form on $\widetilde{X}$, with $k\leq 2p$. Then $\widetilde{f}^{*}\eta$ is $f^{*}\omega$-bounded on $\mathbb{C}^{p}$.
Proof.
Since $\mathbb{C}^{p}$ is simply connected, $f$ lifts to a holomorphic map $\widetilde{f}:\mathbb{C}^{p}\longrightarrow\widetilde{X}$, that is: $\pi\circ\widetilde{f}=f$. In particular, the degeneracy locus of $\widetilde{f}$ is also $\Sigma_{f}$. Let $\eta$ be a $\widetilde{\omega}-$bounded $k$-form on $\widetilde{X}$, with $k\leq 2p$. For any $x\in\mathbb{C}^{p}$, and any tangent vectors $\xi_{1},\cdots,\xi_{k}\in\mathbb{C}^{p}$ at $x$, we have:
@@
Now we can resume the proof of our theorem. $f^{*}\Omega$ defines a positive ($p,p$)-form (hence strongly positive by maximality of the dimension, see Corollary 1.9 in [14], §.1, Chapter III) on $\mathbb{C}^{p}\setminus\Sigma_{f}$. Hence, there exists a smooth positive function $h:\mathbb{C}^{p}\setminus\Sigma_{f}\longrightarrow\mathbb{R}$ such that: $f^{*}\omega_{p}=hf^{*}\Omega$, where $\omega_{p}=\dfrac{\omega^{p}}{p!}$. But since both $f^{*}\omega_{p}$ and $f^{*}\Omega$ are smooth, $h$ extends smoothly to the whole $\mathbb{C}^{p}$. In particular, we have for any $z\in\mathbb{C}^{p}$:
$\sqrt{\binom{n}{p}}=\lvert\omega_{p}(f(z))\rvert_{\omega}=\lvert f^{*}\omega_{p}(z)\rvert_{f^{*}\omega}=h(z)\lvert f^{*}\Omega(z)\rvert_{f^{*}\omega}=h(z)\lvert\Omega(f(z))\rvert_{\omega}$
Thus: $h(z)=\displaystyle\sqrt{\binom{n}{p}}\lvert\Omega(f(z))\rvert^{-1}_{\omega}$, $\forall z\in\mathbb{C}^{p}$. On the other hand, since $\Omega(x)\neq 0$ for any $x\in X$ by positivity, the compactness of $X$ implies that there exists a constant $A>0$ such that: $\dfrac{1}{A}\leq\lvert\Omega(x)\rvert_{\omega}\leq A$. Hence, we get for any $r>0$:
$\text{Vol}_{\omega,f}(\mathbb{B}_{r})=\int_{\mathbb{B}_{r}}f^{*}\omega_{p}=\int_{\mathbb{B}_{r}}hf^{*}\Omega\leq A\sqrt{\binom{n}{p}}\int_{\mathbb{B}_{r}}f^{*}\Omega$ (3.14)
Moreover, we have:
$f^{*}\Omega=\widetilde{f}^{*}\widetilde{\Omega}=\widetilde{f}(\partial\alpha+\bar{\partial}\gamma)\overset{(a)}{=}\widetilde{f}(d(\alpha+\gamma))=d\widetilde{f}(\alpha+\gamma)$ (3.15)
where ($a$) is a consequence of the bidegree constraints. Substituting (3.15) into (3.14) gives:
$\text{Vol}_{\omega,f}(\mathbb{B}_{r})\leq A\sqrt{\binom{n}{p}}\int_{\mathbb{B}_{r}}d\widetilde{f}^{*}(\alpha+\gamma)=A\sqrt{\binom{n}{p}}\int_{\mathbb{S}_{r}}\widetilde{f}^{*}(\alpha+\gamma)\overset{(b)}{\leq}BA_{\omega,f}(\mathbb{S}_{r})\hskip 2.0pt,$ (3.16)
where ($b$) follows from Observation 3.25 (here $\eta=\alpha+\gamma$). On the other hand, the Cauchy-Schwarz inequality yields: